# Efficient Vision-Language-Action Models for Embodied Manipulation: A Systematic Survey

### v2

@@ 2 Foundations and Evolution of VLA Models
Stage II: Introduction of VLMs. A turning point comes with the integration of pre-trained vision-language models (VLMs). RT-2[7] formally introduces the concept of VLA and, for the first time, employs a pre-trained VLM as the backbone. By fine-tuning the VLM on large-scale robot trajectory datasets and converting robot control signals into discrete action tokens, RT-2 successfully transfers the general visual and semantic knowledge acquired from internet data to robot control tasks, achieving a significant leap in generalization. This “VLM-as-a-policy” paradigm quickly becomes mainstream. The release of OpenVLA[8] further accelerates adoption by providing a standardized architecture: strong vision encoders such as SigLIP[9] or DINO-v2[10] for visual feature extraction, a LLaMA tokenizer[11] for text instruction embedding, and a large language model such as LLaMA-7B for high-level reasoning. The LLM output is then used to predict discrete action tokens, which share the same vocabulary as language tokens. Building on this foundation, many follow-up works appear: Octo[12] explores diffusion models as policy heads for continuous action generation; the GR series [13, 14, 15] and VPP[16] adopt ideas from video generation to design pre-training tasks for VLA optimization; ConRFT[17] investigates training paradigms that combine offline and online reinforcement learning. Meanwhile, the release of large and diverse real-world robot datasets such as Open X-Embodiment (OXE)[18] and DROID[19] provides the data support needed to train increasingly large models.
@@ 7 Future Prospects
The preceding sections have systematically reviewed research on optimizing the efficiency of VLA models across their architecture, percepttion feature, action generation, and training and inference pipeline. While significant progress has been made in reducing the computational and resource demands of these models, the path to truly efficient, general, and reliable embodied intelligence is not a simple extrapolation of current techniques. When faced with the complexity, dynamism, and uncertainty of the real physical world, contemporary VLA models still reveal fundamental limitations in their data dependency, perceptual dimensionality, decision-making depth, and capacity for autonomous learning.
This section aims to move beyond a mere inventory of specific technologies to provide a more macroscopic analysis of the core issues shaping the future of the VLA field. We will explore the intrinsic factors that define the trade-offs between model capability and efficiency. This discussion serves not only as a forecast of future research directions but also as an agenda-setting effort to build a solid scientific foundation for the field and advance it toward engineering-grade applications. Our analysis will unfold across five key dimensions: the co-evolution of models and data, breakthroughs in spatio-temporal perception, the cognitive upgrading of action generation, the paradigm shift from imitation to autonomy, and the foundational role of a unified scientific evaluation framework.
7.1 Model and Data: from Scale-Driven Progress to Co-Evolution
Data fuels model intelligence, while models unlock its latent potential. In the VLA domain, the relationship between these two components is far more intricate than the simple heuristic of "more data, bigger models." Future progress hinges on cultivating an ecosystem where model capabilities and data structures evolve in synergy.
A critical starting point is to address the current data predicament in the VLA field. The core issue is not a lack of data in absolute terms, but a systematic scarcity of high-quality, high-density physical interaction data. A hierarchical data pyramid offers a strategic blueprint for the future. At the base lies vast internet-scale data (e.g., web pages, videos), which provides models with broad semantic grounding but lacks an understanding of real-world causality and interaction dynamics. The middle layer consists of simulation data, offering a low-cost, safe environment for trial-and-error learning of complex skills, making it ideal for reinforcement learning. However, its value is perpetually constrained by the Sim-to-Real gap, a discrepancy that extends beyond visual rendering to fundamental differences in physics and contact dynamics. At the apex is the most precious resource: real-world robot data. As the bridge to reality, it provides essential physical feedback. Its high acquisition cost, however, dictates that its primary role should be for validation and fine-tuning, rather than for learning from scratch. A key challenge is to design efficient data flow strategies that enable knowledge transfer from the lower to the upper tiers, thereby maximizing the unique value of each data type.
Building on this, the exploration of Scaling Laws [95] for VLA models must also extend beyond the traditional paradigms established for Large Language Models. The performance metric for VLA systems is not a smooth cross-entropy loss but often a discrete and sparse task success rate. This suggests that VLA scaling behavior may manifest as step-function-like emergent abilities rather than a smooth curve. Furthermore, the independent variables are more complex, encompassing not only model parameters and data volume but also crucial dimensions like modality richness, task diversity, and interaction quality. A pivotal open question remains: does performance improvement stem more from uniform scaling of data volume or from the targeted injection of high-quality interaction data?
This landscape presents a fundamental trade-off between efficiency constraints and the capability ceiling of a model. The current intense focus on efficiency, often realized through aggressive model compression techniques like quantization and pruning, may prove to be a short-sighted strategy. Consequently, future research must shift from merely compressing models to optimizing the allocation of resources. The objective is to systematically identify the Pareto frontier among model architecture, data scale, and task performance within a given hardware budget (e.g., an onboard robotic computing platform). Achieving this balance would enable the co-evolution of efficiency and capability.
7.2 Spatio-Temporal Perception: from 2D Moments to 3D World Model
Current VLA models are constrained by two fundamental perceptual limitations: a spatial limitation due to their reliance on 2D images, and a temporal myopia rooted in the Markovian assumption [96] that decisions depend only on the current frame. These constraints reduce the model to a narrow, short-sighted observer, incapable of forming a coherent and persistent understanding of its three-dimensional environment over time. This, in turn, restricts performance in fine manipulation and long-horizon tasks.
The transition from 2D to 3D perception is not simply a matter of adding another data dimension, but of bridging a representational gap. A 2D image is a view-dependent projection, whereas 3D representations (e.g., point clouds, voxel grids) more closely approximate the physical structure of the world. Understanding in 3D allows reasoning about occlusion, spatial distance, volumetric properties, and object topology, prerequisites for tasks such as threading a needle or assembling complex parts. Yet the computational and memory overhead of 3D data is substantial. Consequently, research is shifting from asking whether 3D is necessary to determining how it can be represented and utilized efficiently. Candidate solutions include voxel-based, point-based, implicit neural, and hybrid methods. Among these, hybrid representations appear especially promising: using high-fidelity, explicit encodings for the robot’s immediate interaction space while employing lightweight, implicit representations for the background.
Beyond spatial perception, overcoming the limitation of instantaneous decision-making requires models to construct a dynamic and persistent world model. This calls for a hierarchical memory system: short-term working memory to track immediate object states, and long-term episodic memory to store extended interaction histories and environmental layouts. The former can be implemented through recurrent networks or Transformer states, while the latter may require more sophisticated, queryable external memory modules.
A central challenge that unifies both spatial and temporal extensions is the trade-off between information fidelity and efficiency. A model cannot store all historical sensory data verbatim; instead, it must perform lossy yet effective compression, learning to discern what information is critical to retain and what can be safely discarded. Designing this dynamic mechanism of memory and forgetting remains an open research problem at the intersection of cognitive science and computer science.
7.3 Action Generation: from Conditioned Reflexes to Deliberative Reasoning
Action is the ultimate output of a VLA system, and the mechanism by which actions are generated directly reflects the system’s level of embodied intelligence. Today many VLA models directly output low-level control signals such as 7-DoF end-effector poses. This approach is a computationally efficient, reactive strategy, but it suffers from limited generalizability, poor interpretability, and an inability to handle complex, multi-step tasks. To move beyond tool-like reflexes toward collaborative agents, action generation must evolve from purely reactive control to structured, deliberative reasoning.
A useful organizing principle is a hierarchy of action abstraction. At the bottom sits low-level control, responsible for immediate, time-critical motor commands and fine adjustments. The middle tier comprises skill primitives: reusable, parameterized behaviors that transform continuous control into a discrete selection and parameterization problem. At the top lies goal-oriented task planning, where long-horizon instructions are decomposed into sequences of skill invocations and subgoals. Structuring action generation along these levels makes complex tasks tractable by delegating routine timing and stability to low levels while reserving higher levels for abstract sequencing and strategy.
Reasoning is central across this hierarchy and can take different forms. Implicit reasoning is embedded inside a single end-to-end model: the system produces actions directly, with reasoning distributed across learned weights. This is often computationally efficient but opaque. Explicit reasoning first constructs an intermediate plan or textual chain-of-thought that is then executed; explicit plans improve interpretability and debuggability but introduce additional computational cost and latency. Each mode has complementary strengths: implicit routes favor speed and fluidity, while explicit routes favor verifiability and corrective intervention.
This dichotomy leads to a fundamental trade-off between reaction speed and decision depth. Excessive deliberation can delay action when timeliness matters; pure reactivity can fail on novel or long-horizon tasks. A pragmatic future direction is a hybrid architecture in which a lightweight reactive module handles routine, low-risk situations, and a heavier explicit reasoning module is invoked selectively for novelty, anomalies, or high-level planning. Implementing such hybridism requires an effective arbitration mechanism that decides when to switch, based on signals like novelty detection, uncertainty estimates, task complexity, or a cost-benefit assessment of computation versus expected performance gain.
7.4 Learning Paradigm: from Imitation to Reinforcement Learning
Imitation Learning (IL), with behavior cloning [97] as a common form, remains the dominant training paradigm for VLA models today, valued for its efficiency and stability. However, its reliance on a fixed dataset limits the scope of learned behaviors to the coverage and quality of demonstrations, making it difficult to generalize to novel or out-of-distribution scenarios. To enable exploration, adaptation, and autonomous improvement, incorporating Reinforcement Learning Fine-Tuning (RLFT) [98] is a promising direction.
Applying RLFT in physical environments presents significant challenges. Sample inefficiency leads to extremely long training times, while safety constraints make unconstrained exploration risky and potentially damaging to both the robot and its surroundings. Reward design for complex, multi-step tasks remains a bottleneck, as it is difficult to specify dense and informative reward signals. In addition, sim-to-real transfer is far from solved, policies optimized in simulation often degrade when deployed in the real world. Thus, the future does not lie in replacing IL with RLFT, but rather in developing safe and efficient strategies that leverage RLFT to address IL’s inherent limitations.
Several recent techniques illustrate possible solutions. Offline RL learns from pre-collected datasets without additional interaction, improving safety and efficiency, and allows models to refine policies using data from demonstrations or previous experiments. Model-based RL improves sample efficiency by first constructing a predictive world model, which can simulate possible outcomes and guide policy refinement without costly real-world trials. Inverse Reinforcement Learning infers a reward function from expert demonstrations, enabling agents to generalize behaviors beyond direct imitation. Together, these approaches provide complementary tools that help VLA models move beyond the limitations of pure imitation, combining safety, efficiency, and adaptability.
We propose a progressive learning ladder for future VLA training. The process begins with IL, giving the model a foundation of basic skills and an initial policy. This is followed by Offline RL, which improves policy robustness without physical risk. Finally, the model undergoes constrained and safety-aware online fine-tuning, adapting to its target environment while incorporating techniques such as RLHF or curriculum learning to further align behavior with task requirements. This structured pipeline combines the stability of imitation with the adaptability of reinforcement learning, offering a practical pathway toward more generalized embodied intelligence.
@@ 7.4 Learning Paradigm: from Imitation to Reinforcement Learning
7.5 Evaluation Criteria: from Fragmented Tests to Unified Evaluation
Progress across VLA research directions, including scaling laws, 3D perception, deliberative reasoning, and RLFT, ultimately relies on a shared foundation: a scientific, unified, and reproducible evaluation framework. The rapid growth of VLA research has highlighted an underlying problem of fragmented evaluation standards and lack of transparency in experimental design. Without common metrics, claims of improved efficiency or generalization are hard to verify.
We advocate for the development of a comprehensive and open research infrastructure for the VLA field. This is a prerequisite for moving from research prototypes to practical deployment. A three-dimensional evaluation framework centered on Efficiency, Performance, and Interpretability provides a clear structure for standardized reporting:
Resource and Efficiency: report model size, number of parameters, inference latency, and hardware used. Include estimates of training time, energy consumption, and memory footprint. This ensures fair comparison across different platforms and scales.
Performance and Robustness: measure task success rates, long-horizon task continuity, and the system’s ability to recover from failure. Evaluate behavior under out-of-distribution conditions, varying environmental factors, or perturbations to ensure reliable real-world deployment.
Interpretability and Auditability: assess how understandable and traceable the model’s decisions are. Include human evaluation scores, mechanisms for decision attribution, visualization of action rationale, and logging of all actions for post-hoc analysis. This is critical for safety and trust in human-robot collaboration.

### v3

@@ 2 Foundations and Evolution of VLA Models
Stage II: Introduction of VLMs. A turning point comes with the integration of pre-trained vision-language models (VLMs). RT-2[7] formally introduces the concept of VLA and, for the first time, employs a pre-trained VLM as the backbone. By fine-tuning the VLM on large-scale robot trajectory datasets and converting robot control signals into discrete action tokens, RT-2 successfully transfers the general visual and semantic knowledge acquired from internet data to robot control tasks, achieving a significant leap in generalization. This "VLM-as-a-policy" paradigm quickly becomes mainstream. The release of OpenVLA[8] further accelerates adoption by providing a standardized architecture: strong vision encoders such as SigLIP[9] or DINO-v2[10] for visual feature extraction, a LLaMA tokenizer[11] for text instruction embedding, and a large language model such as LLaMA-7B for high-level reasoning. The LLM output is then used to predict discrete action tokens, which share the same vocabulary as language tokens. Building on this foundation, many follow-up works appear: Octo[12] explores diffusion models as policy heads for continuous action generation; the GR series [13, 14, 15] and VPP[16] adopt ideas from video generation to design pre-training tasks for VLA optimization; ConRFT[17] investigates training paradigms that combine offline and online reinforcement learning. Meanwhile, the release of large and diverse real-world robot datasets such as Open X-Embodiment (OXE)[18] and DROID[19] provides the data support needed to train increasingly large models.
@@ 7 Future Prospects
The preceding sections have systematically reviewed efforts to improve the efficiency of VLA models across multiple facets, including model architecture, perception feature, action generation, and training-inference paradigms. Despite substantial advances, current models remain constrained when confronted with the complexity, variability, and uncertainty of real-world embodied tasks. Key limitations persist in areas such as redundant data usage, high-dimensional sensory processing, fragmented action sequences, and inefficient learning strategies, all of which impose significant computational and deployment costs.
This section aims to provide a forward-looking perspective, identifying the principal factors that will shape the next generation of efficient VLA systems. Rather than cataloging specific techniques, we examine the trade-offs between capability and efficiency, highlighting where improvements can most effectively reduce computational demand while preserving or enhancing task performance. Our discussion is organized around five interrelated dimensions: the co-optimization of models and data, the development of efficient spatio-temporal perception, the design of compact and continuous action representations, the transition from imitation learning to resource-aware reinforcement adaptation, and the establishment of an efficiency-centered evaluation framework. Together, these dimensions define a roadmap for research that systematically balances capability and efficiency, guiding the field toward practical, generalizable, and resource-conscious embodied intelligence.
7.1 Model and Data: from Scale-Centric Design to Co-Optimized Efficiency
In the evolution of VLA models, the historical paradigm of "larger models trained on more data" is approaching its saturation point. Future progress will increasingly rely on the co-evolution between model and data, where the representational capacity of the model and the structural composition of its training data are jointly optimized. This shift recognizes that efficiency and generalization no longer hinge solely on parameter scaling, but on how effectively a model can exploit, filter, and transfer information across diverse sources of visual, linguistic, and embodied datasets.
However, this trend toward model–data co-design also exposes a new layer of efficiency bottlenecks. The central challenge is no longer data scarcity per se, but data inefficiency, an imbalance between the exponential growth of data volume and the limited computational or memory budgets of embodied agents. Current datasets form a hierarchical "data pyramid", spanning internet-scale corpora, simulation trajectories, and real-world robot interactions. While large-scale data provides rich semantic grounding, the uncurated inclusion of redundant or low-value samples introduces unnecessary computation, prolongs training time, and dilutes high-quality physical feedback. Furthermore, the Sim-to-Real gap compounds this inefficiency: excessive simulation data can obscure the limited but critical contribution of real-world interaction samples. As a result, the overall learning process suffers from diminishing returns, each additional unit of data yields progressively smaller gains relative to its computational cost.
To address these inefficiencies, future efficiency-oriented research should extend beyond model compression toward the joint optimization of model and data. Promising directions include: developing data-centric efficiency frameworks that quantify the marginal utility of different data modalities and sources; exploring selective or curriculum-based data pipelines that dynamically adjust sampling frequency across the data pyramid; investigating joint scaling laws[95] that model how performance scales not only with parameters and data size but also with data quality and diversity. Such approaches would enable VLA models to allocate their limited training and inference budgets more intelligently, achieving a Pareto-optimal balance between data volume, model complexity, and embodied task performance. Ultimately, this co-evolutionary perspective reframes efficiency not as a constraint but as a guiding principle for sustainable scaling in embodied intelligence.
7.2 Spatio-Temporal Perception: from 2D Frames to Efficient 3D Representations
Current VLA models are constrained by two fundamental perceptual limitations: a spatial limitation rooted in the predominance of 2D image inputs, and a temporal myopia stemming from the Markovian assumption [96] that decisions rely solely on the current frame. As embodied tasks grow in spatial and temporal complexity, requiring reasoning about occlusion, volumetric relations, and long-horizon intent, these simplifications increasingly fail. The natural evolution is therefore toward richer, three-dimensional and temporally grounded perceptual systems: explicit or implicit 3D representations (point clouds, voxel grids, neural fields), multi-view fusion, and persistent memory structures that record and summarize interaction histories.
Moving to such spatio-temporal representations, however, raises a practical efficiency dilemma. The added spatial fidelity and temporal context multiply the number of perceptual tokens (3D points, multi-view patches, or frame-wise embeddings) fed into the downstream LLM backbone. In VLA architectures where LLM perform the multimodal reasoning and action decoding, this token explosion translates directly into computational increases in attention and feedforward costs, larger activation footprints during inference, and greater latency, all of which threaten real-time control. Moreover, unselective retention of past observations blurs signal-to-noise ratios: vast swathes of redundant or low-value sensory history consume compute without commensurate gains in task success.
Given these constraints, future research should prioritize principled strategies to preserve spatial and temporal understanding while containing the token burden. At the spatial level, work may investigate task-aware 3D summarization that encodes high-fidelity geometry only in regions relevant to imminent interaction and represents background context with compact implicit descriptors. Temporally, approaches that combine short-term dense state tracking with long-term sparse summaries (keyframes or episodic sketches) can maintain continuity without unbounded history growth. Across modalities, semantic-guided filtering, using language or task priors to gate perception, and learnable token pruning during fusion can remove redundant signals before they enter the costly reasoning stage. Finally, evaluations should explicitly report the downstream backbone cost (token counts, attention FLOPs, inference latency) alongside success rates, to reveal practical Pareto frontiers between representational richness and deployable efficiency.
7.3 Action Generation: from Chunked Commands to Compact Continuous Control
Action generation stands at the core of a VLA system’s embodiment capability. Current models predominantly operate under a chunked-control paradigm, where the model outputs discrete segments of action chunks to guide robot motion. This design enables efficient low-frequency control suitable for pick-and-place tasks. However, as robotic systems advance toward high-precision and high-frequency operations, such as assembly or dexterous manipulation, the temporal granularity and dynamic continuity required for control will inevitably increase. The emerging trend is thus a shift from low-rate, discrete command generation toward continuous and context-aware control, where the model not only decides what to do but also how to modulate its actions in real time.
This evolution brings a new class of efficiency bottlenecks. Increasing action chunk length or control frequency directly expands the number of output tokens, shifting the computational burden from the perception side to the language backbone itself. The autoregressive decoding cost grows nonlinearly with sequence length, making real-time inference on embodied platforms increasingly challenging. Moreover, maintaining temporal consistency across chunks becomes difficult, because the lack of smooth transition between successive output segments often causes discontinuous motion or instability in execution. Attempts to embed explicit reasoning, such as Chain-of-Thought planning, further exacerbate latency and memory demands, creating an acute tension between deliberation depth and control responsiveness. In embodied settings where control cycles operate at tens or hundreds of hertz, such reasoning overhead is often untenable.
To reconcile these opposing demands, future research must explore compact and hierarchical representations of actions that balance expressiveness and efficiency. One promising direction is to organize control outputs into multi-level abstractions: low-level motion primitives handled reactively, and higher-level intentions encoded through latent or parameterized tokens, thereby compressing output bandwidth without sacrificing control fidelity. Another key avenue is developing mechanisms for cross-chunk temporal coherence, for instance, caching or reusing hidden states between consecutive action segments to ensure smooth dynamics while avoiding redundant computation. Finally, the integration of reasoning within VLA must evolve toward reactive reasoning: lightweight, on-demand internal planning that preserves interpretability without incurring prohibitive latency. Such designs point toward a future where VLA systems achieve a true synthesis of reasoning and control, thinking and acting as one coherent, efficient process.
7.4 Learning Paradigms: from Imitation to Efficient Reinforcement Adaptation
Imitation Learning (IL), typically instantiated as behavior cloning [97], remains the dominant paradigm for training VLA models due to its simplicity, stability, and high sample efficiency. By leveraging expert demonstrations, IL allows models to learn robust visuomotor mappings without costly trial-and-error interaction. However, this paradigm fundamentally constrains the ceiling of performance: the learned policy cannot exceed the coverage and quality of the demonstration dataset. As VLA systems pursue greater autonomy, adaptability, and generalization, the field is gradually shifting toward hybrid paradigms that incorporate elements of Reinforcement Learning (RL), enabling models not only to imitate but also to explore, optimize, and refine their own behavior beyond expert priors.
Yet this transition introduces a new spectrum of efficiency challenges. RL in embodied environments is notoriously sample-inefficient, requiring millions of interactions to achieve convergence, and such extensive exploration is impractical or unsafe on physical robots. Furthermore, VLA models compound the difficulty due to their large multimodal backbones, which increase the cost of each policy update. Reward specification for complex, language-conditioned manipulation tasks remains an open bottleneck, often producing sparse or ambiguous learning signals. Sim-to-real transfer further amplifies inefficiency: even policies refined in high-fidelity simulators tend to degrade in real-world deployment due to mismatched dynamics and sensing noise. Thus, while RL holds the promise of autonomy, its naive application to embodied agents risks becoming prohibitively expensive in both data and computation.
A promising research direction lies in redefining RL through the lens of efficiency. Rather than replacing imitation learning, future systems may employ progressive training pipelines that combine the strengths of both paradigms. An efficient learning hierarchy could begin with imitation learning to initialize a stable policy, followed by large-scale offline reinforcement fine-tuning using pre-collected interaction data to enhance generalization, and finally a limited stage of safety-constrained online adaptation in real environments. Complementary efficiency techniques, such as model-based rollouts, experience replay across multimodal contexts, and adaptive reward shaping, may further improve sample utilization and reduce hardware wear. In this view, the future of VLA training lies not in maximizing exploration, but in optimally allocating interaction budgets, achieving autonomous learning under strict efficiency constraints.
@@ 7.4 Learning Paradigms: from Imitation to Efficient Reinforcement Adaptation
7.5 Evaluation Criteria: from Fragmented Measures to Efficiency-Centric Benchmarking
A shared and standardized evaluation framework is essential for meaningful progress in VLA research, particularly when efficiency is a central concern. Current VLA studies report heterogeneous metrics, rely on diverse datasets, and employ varied hardware, making cross-comparison difficult and obscuring true efficiency gains. Without a rigorous, reproducible benchmark, it is nearly impossible to quantify whether proposed techniques actually reduce computational cost without compromising task performance.
We argue that a structured, open benchmarking ecosystem can directly drive efficiency-oriented research. A three-dimensional evaluation framework, centered on resource efficiency, task performance, and interpretability, can guide the design and comparison of models with a dual focus on capability and computational cost:
Resource and Efficiency: Standardize reporting of model size, parameter count, inference latency, and hardware setup, alongside training time, memory footprint, and energy consumption. Such transparency allows equitable comparisons across different scales and platforms, and highlights the trade-offs between computational demand and task performance.
Performance and Robustness: Beyond task success rates, evaluation should consider long-horizon stability, recovery from failure, and resilience to environmental variations or sensor noise. Testing under distribution shifts and real-world perturbations ensures that performance metrics reflect true embodied competence rather than overfitting to narrow benchmarks.
Interpretability and Auditability: As embodied systems increasingly share control spaces with humans, understanding why a model acts is as important as what it does. Evaluation should therefore incorporate human-judged interpretability, visualization of action rationales, decision attribution mechanisms, and transparent action logging for post-hoc analysis—foundations for safety, accountability, and trust.